OpenStudy (peterg2001):

help will medal Solve the following equation, and check your solution. Be sure to show all work in order to receive full credit. Solve -17 + n/5 = 33.

OpenStudy (anonymous):

-17 + n/5 = 33 +17 +17 n/5 = 50 n/5 x 5 = 50 x 5 n = 250 :)

OpenStudy (anonymous):

n=250